OpenStudy (chrisplusian):

Find parametric equations and symmetric equations for the line (use parameter t). The line through (1,-4,4) an parallel to the line: x+2=y/2=z-5

OpenStudy (chrisplusian):

I was thinking that snce it says that x+2=y/2=z-5 that these are=t? Then couldn't I set each prtion equal to t? t=x+2,t=y/2, t=z-5? Then if I solve for x, y,and z wouldn't that be the parametric equations?